A project pays off $ 2,160 right away, has $ 5,065 in costs at the end of the second year and again pays off 3,961 at the end of the third year. If MARR is 16 %, what is the project's approximate ERR? (Note: Give your answer in percentages without % sign and keep two decimal places)
Answer #1

PW of project outflows = 5065 * (P/F,16%,2)

= 5065 *0.743163

= 3764.12

FW of project inflows = 2160 *(F/P,16%,3) + 3961

= 2160 *1.560896 + 3961

= 7332.54

Let ERR be i%, then

3764.12*(1+i)^3 = 7332.54

(1+i)^3 = 7332.54 / 3764.12 = 1.948007556

1+i = 1.948007556^(1/3) = 1.2489

i = 1.2489 - 1 = 0.2489 ~ 24.89%